By doubling each dimension thearea of a triangle increased from 6 square miles to 24 square miles. Find the percent increase in area.
Found 2 solutions by JBarnum, amoresroy:
Answer by JBarnum(2146)

(24)/6=4 4times 100%=400%
so 400% increase in area
6=100%
12=200%
18=300%
24=400%

Answer by amoresroy(361)

% increase in area = (24-6)/6 * 100%
= 18/6 * 100%
= 300%